joye built a rectangular gared whose lenghts was eight feet more than triple it's width. if the perimeter was two hundred sixteen feet, what is the length of the garden
Answer by oberobic(2304)

P = perimeter
L = length
W = width
.
P = 2L + 2W = 2(L+W)
.
P = 216 ft. (given)
.
The length is 8 ft more than 3 times the width.
L = 3*W + 8
.
2(L+W) = 216
.
Divide both sides by 2
.
L+W = 108
.
Substitute for L
.
3*W+8 + W = 108
.
4W + 8 = 108
4W = 100
W = 25
.
The width is 25 ft.
.
Substitute to find L.
.
L+W = 108
L + 25 = 108
L = 83
.
The length is 83 ft.
.
Always check your work!
.
In this case, use the other statement and equation.
The length is 8 ft more than 3 times the width.
L = 3W + 8
L = 83
3W + 8 = 3(25) + 8
3*25 + 8 = 83
Correct!
.
Re-read the question to make sure you answer it.
.
What is the length of the garden?
The garden is 83 ft. in length.
.
Done.